OpenStudy (anonymous):

What is the solution of the system? Use elimination. 3x – 4y = –24 x + y = –1 A. (–4, 3) B. (0, 6) C. (3, 4) D. (4, 3)

OpenStudy (anonymous):

@BacardiDavis D?!?

OpenStudy (jack1):

you have x + y = –1 (EQN 1) and you want to use elimination... so 3 (EQN 1) = 3x + 3y = -3 so 3x = -3 -3y

OpenStudy (anonymous):

so A? lol

OpenStudy (jack1):

call 3x = -3 -3y (eqn 1a) so 3x = 4y - 24 and 3x = -3 -3y so 4y - 24 = -3 -3y

OpenStudy (jack1):

can you get it from here?

OpenStudy (anonymous):

yeah thanks :p

OpenStudy (anonymous):

:p